No relevant resource is found in the selected language.

This site uses cookies. By continuing to browse the site you are agreeing to our use of cookies. Read our privacy policy>

Reminder

To have a better experience, please upgrade your IE browser.

upgrade

Solution for alarms CPC_RSVP_NB_DOWN & CPW_OMS_TEL_DEG inside ASON Network

Publication Date:  2017-10-14 Views:  56 Downloads:  0
Issue Description

1. Inside link ASON of SPAN "MEDELLIN" to "MANIZALES" reported alarms CPC_RSVP_NB_DOWN & CPW_OMS_TEL_DEG:

 

 

 

2. Alarms present in different node as LA PINTADA, SALAMINA, MANIZALES: 

 

 

3.  the communication status has: Receiving Failed of channels: D1-D3, D4-D12, FE_DCN of board ST2 port RM2/TM2

Alarm Information

The CPC_RSVP_NB_DOWN alarm indicates that the communication between neighboring NEs fails, and the RSVP message from the neighboring NE cannot be received. The alarm is cleared when the RSVP communication between the neighboring NEs is restored, and an event indicating the end of this alarm is reported. In addition, the alarm is not reported when the neighboring NE does not exist. Impact on the System: The service signals are not affected, but the normal signaling process of the service is affected. In this case, certain operations cannot be performed. For example, service optimization or rerouting on the two NEs fails.



The CPW_OMS_TEL_DEG alarm indicates that the OMS TE link is degraded. Impact on the System: The service is not affected; however, the creation of a service, optimize the existing service, or rerouting of the existing service on the TE link fails.



Handling Process

1. check diagram and obtain Data collector file of sites in order to analyze logs

2. Go to TE link, do synchronize, but still not ok. 

3. Check TE link status and found one site have no any TE link active with other equipment. Check control link, the same.
 
4. Then check the ASON Domain, found this site are separated from others.

Root Cause

apply cold reset to the EFI board on site SALAMINA

Solution

【Resolution Summary】: apply cold reset to EFi board in order to resolve it two alarms
【Resolution Details】Solution Analysis:the dcc is all rx_f state, we think the problem is on the EFI board, after cold reset the EFi board, the dcc is normal now, and all the alarm cleared

Resolution Detail: when apply cold reset to the EFi board, the dcc is normal now, because the board restore configuration and all the alarm cleared . please replace this EFI if same issue appear again.

 

 //Risk of Solution:Non-Risk Operation

END